# Nearly-Linear Time Seeded Extractors with Short Seeds

### v1

@@ 1.3 Technical Overview
In a nutshell, we obtain 1 by following two standard high-level steps:
@@
In order to implement Item 1, we need to use fast condensers with short seeds. Luckily for us, some existing state-of-the-art constructions of condensers already satisfy this property, although, to the best of our knowledge, this has not been observed before. We argue this carefully in Section 3.3.
For example, the “lossy Reed-Solomon condenser” from [GUV09] interprets the source as a polynomial $f\in\mathds{F}_{q}[x]$ of degree $d\leq n/\log q$ and the seed $y$ as an element of $\mathds{F}_{q}$, and outputs $\mathsf{RSCond}(f,y)=(f(y),f(\zeta y),\dots,f(\zeta^{m^{\prime}}y))$, for an appropriate $m^{\prime}$ and field size $q$, with $\zeta$ a primitive element of $\mathds{F}_{q}$. Evaluating $\mathsf{RSCond}(f,y)$ corresponds to evaluating the same polynomial $f$ on multiple points in $\mathds{F}_{q}$. This is an instance of the classical problem of multipoint evaluation in computational algebra, for which we know fast and practical algorithms (e.g., see [vzGG13, Chapter 10] or Lemma 2.1) running in time $\widetilde{O}((d+m^{\prime})\log q)=\widetilde{O}(n)$, since $d\leq n/\log q$ and if $m^{\prime}\leq n/\log q$.
A downside of this condenser is that it requires knowing a primitive element $\zeta$ of $\mathds{F}_{q}$ with $q=\operatorname{poly}(n/\varepsilon)$. As discussed above, if we know the factorization of $q-1$ and $q$ is a power of $2$, then we can find such a primitive element in time $\operatorname{polylog}(q)$. Beyond that, having access to such primitive elements, which only need to be computed once independently of the source and seed, is reasonable in practice. Therefore, we may leave this as a one-time preprocessing step.
The lossless “KT condenser” from [KT22] has a similar flavor. It interprets the source as a polynomial $f\in\mathds{F}_{q}[x]$ and the seed $y$ as an evaluation point, and outputs $\mathsf{KTCond}(f,y)=(f(y),f^{\prime}(y),\dots,f^{(m^{\prime})}(y))$, for some appropriate $m^{\prime}$. The problem of evaluating several derivatives of the same polynomial $f$ on the same point $y$ (sometimes referred to as Hermite evaluation) is closely related to the multipoint evaluation problem above, and can also be solved in time $\widetilde{O}(n)$. [^6] Interestingly, recent works used other useful computational properties of the KT condenser. Cheng and Wu [CW24] crucially use the fact that the KT condenser can be computed in NC1. Doron and Tell [DT23] use the fact that the KT condenser is logspace computable for applications in space-bounded derandomization. Evaluating the KT condenser does not require preprocessing. On the other hand, it only works when the min-entropy $k\geq C\log^{2}(n/\varepsilon)$ for a large constant $C>0$, where $n$ is the source length and $\varepsilon$ the target error of the condenser.
@@
We have seen that there are fast condensers with short seeds. It remains to realize Item 2. Because of the initial condensing step, we may essentially assume that our $n$-bit weak source $X$ has min-entropy $k\geq(1-\delta)n$, for an arbitrarily small constant $\delta>0$. In this case, we would like to realize in time $\widetilde{O}(n)$ and with overall seed length $O(\log(n/\varepsilon))$ what we see as the most natural approach to seeded extraction from high min-entropy sources:
@@
    Use a fresh short seed to transform $X$ into a block source $Z=Z_{1}\circ Z_{2}\circ\cdots\circ Z_{t}$ with geometrically decreasing blocks, where $\circ$ denotes string concatenation. A block source has the property that each block $Z_{i}$ has good min-entropy even conditioned on the values of blocks $Z_{1},\dots,Z_{i-1}$.
@@
    Perform block source extraction on $Z$ using another fresh short seed. Due to its special structure, we can extract a long random string from $Z$ using only the (small) seed length associated with extracting randomness from the smallest block $Z_{t}$, which has length $O(\log(n/\varepsilon))$.
The classical approach to Item 2 where we iteratively apply extractors based on universal hash functions with increasing output lengths to the blocks of $Z$ from right to left is easily seen to run in time $\widetilde{O}(n)$ and requires a seed of length $O(\log(n/\varepsilon))$ if, e.g., we use the practical extractors of [TSSR11, HT16]. Therefore, we only need to worry about realizing Item 1.
A standard approach to Item 1 would be to use an averaging sampler to iteratively sample subsequences of $X$ as the successive blocks of the block source $Z$, following a classical strategy of Nisan and Zuckerman [NZ96] (improved by [RSW06, Vad04]). We do know averaging samplers running in time $\widetilde{O}(n)$ (such as those based on random walks on a carefully chosen expander graph). However, this approach requires a fresh seed of length $\Theta(\log(n/\varepsilon))$ per block of $Z$. Since $Z$ will have roughly $\log n$ blocks, this leads to an overall seed of length $\Theta(\log^{2}n+\log(1/\varepsilon))$, which is too much for us.
Instead, we provide a new analysis of a sampler based on bounded independence, that runs in time $\widetilde{O}(n)$ and only requires a seed of length $O(\log(n/\varepsilon))$ to create the entire desired block source. We give the construction, which may be of independent interest, in Section 3.2. The caveat of this “block source creator” is that it only works as desired when the target error $\varepsilon\geq 2^{-k^{c}}$ for some small constant $c>0$. Combining these realizations of Items 1 and 2 yields the desired $\widetilde{O}(n)$-time extractor with order-optimal seed length $O(\log(n/\varepsilon))$ and output length $(1-\eta)n$ for arbitrary constant $\eta>0$, provided that $\varepsilon\geq 2^{-k^{c}}$. See Theorem 5.1 for the formal statement.
@@
We saw above that combining the ideal approach to seeded extraction from high min-entropy sources with the new analysis of the bounded independence sampler yields a conceptually simple construction with the desired properties when the error is not too small (or alternatively, whenever the entropy guarantee is large enough). However, we would like to have $\widetilde{O}(n)$-time seeded extraction with $O(\log(n/\varepsilon))$ seed length and large output length for all ranges of parameters.
@@ 1.4 Future Work
  • Remove the preprocessing step that our constructions behind 1 require when $k<C\log^{2}(n/\varepsilon)$.
  • On the practical side, develop concrete implementations of seeded extractors with near-optimal seed length and large output length. In particular, we think that our non-recursive construction in Section 5.1 holds promise in this direction.
@@ 2.1 Notation
We often use uppercase roman letters to denote sets and random variables – the distinction will be clear from context. We denote the support of a random variable $X$ by $\mathsf{supp}(X)$, and, for a random variable $X$ and set $S$, we also write $X\sim S$ to mean that $X$ is supported on $S$. For a random variable $X$, we write $x\sim X$ to mean that $x$ is sampled according to the distribution of $X$. We use $U_{d}$ to denote a random variable that is uniformly distributed over $\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$. For two strings $x$ and $y$, we denote their concatenation by $x\circ y$. Given two random variables $X$ and $Y$, we denote their product distribution by $X\times Y$ (i.e., $\Pr[X\times Y=x\circ y]=\Pr[X=x]\cdot\Pr[Y=y]$. Given a positive integer $n$, we write $[n]=\{1,\dots,n\}$. For a prime power $q$, we denote the finite field of order $q$ by $\mathds{F}_{q}$. We denote the base-$2$ logarithm by $\log$.
@@ 2.2 Model of Computation
We work in the standard, multi-tape, Turing machine model with some fixed number of work tapes. In particular, there exists a constant $C$ such that all our claimed time bounds hold whenever we work with at most $C$ work tapes. This also implies that our results hold in the RAM model, wherein each machine word can store integers up to some fixed length, and standard word operations take constant time. In Section 4 we will give, in addition to the standard Turing machine model bounds, an improved runtime bound that is dedicated to the logarithmic-cost RAM model.
2.3 Fast Finite Fields Operations
For a prime power $q=p^{\ell}$, we let $M_{q}(d)$ be the number of field operations required to multiply two univariate polynomials over $\mathds{F}_{q}$ of degree less than $d$, and $M_{q}^{\mathsf{b}}(d)$ be the bit complexity of such a multiplication, so $M_{q}^{\mathsf{b}}(d)\leq M_{q}(d)\cdot T(q)$, where we denote by $T(q)$ an upper bound on the bit complexity of arithmetic operations in $\mathds{F}_{q}$. When $\ell=1$, Harvey and van der Hoeven [HvdH19, HvdH21] showed that
$M^{\mathsf{b}}_{q}(d)=O(d\log q\cdot\log(d\log q)\cdot 4^{\max(0,\log^{*}\!d-\log^{*}\!q)}),$
and in general, $M_{q}(d)=d\cdot\log d\cdot 2^{O(\log^{\star}n)}$ [Für09]. [^8] If $\mathds{F}_{q}$ contains a $d$-th root of unity, one can get $M_{q}(d)=d\log d$ from the classic FFT algorithm [CT65]. For a simpler algorithm attaining the bound $M_{q}(d)=d\log d\operatorname{loglog}d$, see [vzGG13, Sections 8,10]. See also [HvdH22] for a widely-believed conjecture under which $M_{q}(d)=d\log d$ always holds. When $p=2$, we can use Schönhage’s algorithm [Sch77] to get $M_{q}^{\mathsf{b}}(d)=O(d\log d\cdot\operatorname{loglog}d\cdot M_{q}(\log q)),$ where we relied on the fact that addition and multiplication in $\mathds{F}_{q}$ can be done in time $M_{q}(\ell)=O(\ell\cdot\log\ell\cdot\operatorname{loglog}\ell)$. Overall, when $d\leq q\leq 2^{d}$, and $q$ is either a prime or a power of two, $M_{q}^{\mathsf{b}}(d)=d\log d\cdot\widetilde{O}(\log q)$. We will use fast multi-point evaluation and fast computation of derivatives (together with the preceding bounds on $M_{q}^{\mathsf{b}}$).
Lemma 2.1 ([BM74], see also [vzGG13, Chapter 10]).
Let $d\in\mathbb{N}$, and let $q$ be a prime, or a power of $2$. Then, given a polynomial $f\in\mathds{F}_{q}[X]$ of degree at most $d$, the following holds.
@@ 2.3 Fast Finite Fields Operations
    For $t\leq d$ and $\alpha\in\mathds{F}_{q}$, one can compute the derivatives $f(\alpha),f^{\prime}(\alpha),\ldots,f^{(t)}(\alpha)$ in time $O(M_{q}(d)\cdot\log d)=d\log^{2}d\cdot\widetilde{O}(\log q)$.
Note that when $q\leq 2^{d}$, we can bound $O(M_{q}(d)\cdot\log d)$ by $\widetilde{O}(d)\cdot\log q$ [^9] The looser bound of $\widetilde{O}(d)\cdot\log q$, when $q\leq 2^{d}$, is also a bound for an arbitrary $\mathds{F}_{q}$, and can be achieved with simpler algorithms than the ones cited.
For a comprehensive discussion of fast polynomial arithmetic, see Von Zur Gathen and Gerhard’s book [vzGG13] (and the more recent important developments [HvdH21]).
@@ 2.4 Statistical Distance, Entropy
We present some relevant definitions and lemmas about the statistical distance and min-entropy.
@@
$\Delta(X,Y)=\sup_{\mathcal{T}\subseteq\mathcal{S}}|\Pr[X\in\mathcal{T}]-\Pr[Y\in\mathcal{T}]|=\frac{1}{2}\sum_{x\in\mathcal{S}}|\Pr[X=x]-\Pr[Y=x]|.$
@@
Above, and throughout the paper, we use base-$2$ logarithms.
@@
Definition 2.7 (smooth min-entropy).
We say that a random variable $X$ has $\varepsilon$-smooth min-entropy at least $k$, denoted by $\mathbf{H}_{\infty}^{\varepsilon}(X)\geq k$, if there exists a random variable $X^{\prime}$ such that $X\approx_{\varepsilon}X^{\prime}$ and $\mathbf{H}_{\infty}(X^{\prime})\geq k$.
@@ 2.5 Extractors and Condensers
A random variable $X$ is an $((n_{1},n_{2},\dots,n_{t}),(k_{1},\dots,k_{t}))$-block source if we can write $X=X_{1}\circ X_{2}\circ\cdots\circ X_{t}$, each $X_{i}\in\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n_{i}}$, where $\widetilde{\mathbf{H}}_{\infty}(X_{i}|X_{1},\ldots,X_{i-1})\geq k_{i}$ for all $i\in[s]$. In the special case where $k_{i}=\alpha n_{i}$ for all $i\in[t]$, we say that $X$ is an $((n_{1},n_{2},\dots,n_{t}),\alpha)$-block source.
We say that $X$ is an exact block source if $\mathbf{H}_{\infty}(X_{i}|X_{1}=x_{1},\ldots,X_{i-1}=x_{i-1})\geq k_{i}$ for any prefix $x_{1},\dots,x_{i-1}$. Lemma 2.6 tells us that any $((n_{1},\ldots,n_{t}),\alpha)$-block-source is $\varepsilon$-close to an exact $((n_{1},\ldots,n_{t}),(1-\zeta)\alpha)$-block-source, where $\varepsilon=\sum_{i=1}^{t}2^{-\alpha\zeta n_{i}}$.
@@
where $Y$ is uniformly distributed over $\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$ and is independent of $X$ and $U_{m}$ is uniformly distributed over $\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m}$. We say that $\mathsf{Ext}$ is strong if $\mathsf{Ext}(X,Y)\circ Y\approx_{\varepsilon}U_{m+d}$.
@@
$\mathsf{Ext}(X,Y)\circ Y\circ W\approx_{\varepsilon}U_{m+d}\circ W,$
where $Y$ is uniformly distributed over $\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$ and is independent of $X$ and $U_{m+d}$ is uniformly distributed over $\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m+d}$ and independent of $W$.
@@
By Lemma 2.6, every strong $(k,\varepsilon)$-seeded extractor $\mathsf{Ext}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\to\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m}$ is also an average-case strong $(k^{\prime}=k+\log(1/\varepsilon),\varepsilon^{\prime}=2\varepsilon)$-seeded extractor.
@@
A function $\mathsf{Cond}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\rightarrow\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m}$ is a $(k,k^{\prime},\varepsilon)$ (seeded) condenser if the following holds. For every $(n,k)$-source $X$, $\mathbf{H}_{\infty}^{\varepsilon}(\mathsf{Cond}(X,Y))\geq k^{\prime}$, where $Y$ is uniformly distributed over $\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}$ and is independent of $X$.
We say that $\mathsf{Cond}$ is strong if $Y\circ\mathsf{Cond}(X,Y)$ is $\varepsilon$-close to some distribution $Y\circ D$ with min-entropy $k^{\prime}$ (and note that here, necessarily, $d$ bits of entropy come from the seed). Finally, we say that $\mathsf{Cond}$ is lossless if $k^{\prime}=k+d$.
We also define extractors tailored to block sources.
@@
A function $\mathsf{BExt}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n_{1}}\times\cdots\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n_{t}}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\to\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{m}$ is a $(k_{1},\dots,k_{t},\varepsilon)$ strong block-source extractor if for any $((n_{1},n_{2},\dots,n_{t}),(k_{1},\dots,k_{t}))$-block-source $X$,
$\mathsf{BExt}(X,Y)\circ Y\approx_{\varepsilon}U_{m+d},$
@@
We will also require the following extractors based on the leftover hash lemma and fast hash functions. We state a result from [TSSR11] which requires seed length $d\approx 2m$, where $m$ is the output length. It is possible to improve the seed length to $d\approx m$, but this requires the input length $n$ to be structured [HT16].
Lemma 2.14 (fast hash-based extractors [TSSR11, Theorem 10], adapted. See also [HT16, Table I]).
@@
Note that by appending the seed to the output of the extractor, we can get the following: There exists a constant $c$ such that for any constant $\theta\leq\frac{1}{3}$, $d\geq c\log(n/\varepsilon)$ and $k\geq\theta d+c\log(1/\varepsilon)$, there exists a strong $(k,\varepsilon)$-seeded extractor $\mathsf{Ext}\colon\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{d}\rightarrow\mathopen{}\mathclose{{}\left\{{0,1}}\right\}^{(1+\theta)d}$.

### v2

@@ 1.3 Technical Overview
In a nutshell, we obtain ˜1 by following two standard high-level steps:
@@
In order to implement item˜1, we need to use fast condensers with short seeds. Luckily for us, some existing state-of-the-art constructions of condensers can already be computed in nearly-linear time, although, to the best of our knowledge, this has not been observed before. We argue this carefully in section˜3.3.
For example, the “lossy Reed-Solomon condenser” from [GUV09] interprets the source as a polynomial $f\in\mathds{F}_{q}[x]$ of degree $d\leq n/\log q$ and the seed $y$ as an element of $\mathds{F}_{q}$, and outputs $\mathsf{RSCond}(f,y)=(f(y),f(\zeta y),\dots,f(\zeta^{m^{\prime}-1}y))$, for an appropriate $m^{\prime}$ and field size $q$, with $\zeta$ a primitive element of $\mathds{F}_{q}$. Evaluating $\mathsf{RSCond}(f,y)$ corresponds to evaluating the same polynomial $f$ on multiple points in $\mathds{F}_{q}$. This is an instance of the classical problem of multipoint evaluation in computational algebra, for which we know fast and practical algorithms (e.g., see [vG13, Chapter 10] or lemma˜2.1) running in time $\widetilde{O}((d+m^{\prime})\log q)=\widetilde{O}(n)$, since $d\leq n/\log q$, and if $m^{\prime}\leq n/\log q$.
A downside of this condenser is that it requires knowing a primitive element $\zeta$ of $\mathds{F}_{q}$ with $q=\operatorname{poly}(n/\varepsilon)$. But note that finding this primitive element only needs to be done once for a given set of parameters $(n,k,\varepsilon,m)$ independently of the actual seed and input source, and so we leave it as a one-time preprocessing step. As discussed in remark˜3.16, we have the freedom of choosing $q$ to be prime, and in that case we can find this primitive element in randomized time $\operatorname{polylog}(q)=\operatorname{polylog}(n/\varepsilon)$.
The lossless “KT condenser” from [KT22] has a similar flavor. It interprets the source as a polynomial $f\in\mathds{F}_{q}[x]$ and the seed $y$ as an evaluation point, and outputs $\mathsf{KTCond}(f,y)=(f(y),f^{\prime}(y),\dots,f^{(m^{\prime}-1)}(y))$, for some appropriate $m^{\prime}$. The problem of evaluating several derivatives of the same polynomial $f$ on the same point $y$ (sometimes referred to as Hermite evaluation) is closely related to the multipoint evaluation problem above, and can also be solved in time $\widetilde{O}(n)$. [^9] Interestingly, recent works used other useful computational properties of the KT condenser. Cheng and Wu [CW24] crucially use the fact that the KT condenser can be computed in NC1. Doron and Tell [DT23] use the fact that the KT condenser is logspace computable for applications in space-bounded derandomization. Furthermore, evaluating the KT condenser only requires preprocessing when $\varepsilon$ is very small. On the other hand, it only works when the min-entropy $k\geq C\log^{2}(n/\varepsilon)$ for a large constant $C>0$, where $n$ is the source length and $\varepsilon$ the target error of the condenser.
@@
We have seen that there are fast condensers with short seeds. It remains to realize item˜2. Because of the initial condensing step, we may essentially assume that our $n$-bit weak source $X$ has min-entropy $k\geq(1-\delta)n$, for an arbitrarily small constant $\delta>0$. In this case, we would like to realize in time $\widetilde{O}(n)$ and with overall seed length $O(\log(n/\varepsilon))$ what we see as the most natural approach to seeded extraction from high min-entropy sources:
@@
    Use a fresh short seed to transform $X$ into a block source $Z=(Z_{1},Z_{2},\dots,Z_{t})$ with geometrically decreasing blocks. A block source has the property that each block $Z_{i}$ has good min-entropy even conditioned on the values of blocks $Z_{1},\dots,Z_{i-1}$.
@@
    Perform block source extraction on $Z$ using another fresh short seed. Due to its special structure, we can extract a long random string from $Z$ using only the (small) seed length associated with extracting randomness from the smallest block $Z_{t}$.
Similar approaches were taken in [NZ96, ZUC96], but they do not support logarithmic seed and low-error (see section˜4.1). The classical approach to item˜2 where we iteratively apply extractors based on universal hash functions with increasing output lengths to the blocks of $Z$ from right to left is easily seen to run in time $\widetilde{O}(n)$ and requires a seed of length $O(\log(n/\varepsilon))$ if, e.g., we use the practical extractors of [TSS+11, HT16]. Therefore, we only need to worry about realizing item˜1.
A standard approach to item˜1 would be to use an averaging sampler to iteratively sample subsequences of $X$ as the successive blocks of the block source $Z$, following a classical strategy of Nisan and Zuckerman [NZ96] (improved by [RSW06, VAD04]). We do know averaging samplers running in time $\widetilde{O}(n)$ (such as those based on random walks on a carefully chosen expander graph). However, this approach requires a fresh seed of length $\Theta(\log(n/\varepsilon))$ per block of $Z$. Since $Z$ will have roughly $\log n$ blocks, this leads to an overall seed of length $\Theta(\log^{2}n+\log(1/\varepsilon))$, which is too much for us.
Instead, we provide a new analysis of a sampler based on bounded independence, that runs in time $\widetilde{O}(n)$ and only requires a seed of length $O(\log(n/\varepsilon))$ to create the entire desired block source. However, this block source has blocks of increasing lengths, whereas we need decreasing blocks to perform the block source extraction. We remedy that by sub-sampling from each block using a standard expander random walk sampler.
We give the construction, which may be of independent interest, in section˜3.2. The caveat of this construction is that it only works as desired when the target error $\varepsilon\geq 2^{-k^{c}}$ for some small constant $c>0$. See section˜4.1 for the formal analysis.
@@
We saw above that combining the ideal approach to seeded extraction from high min-entropy sources with the new analysis of the bounded independence sampler yields a conceptually simple construction with the desired properties when the error is not too small. However, we would like to have $\widetilde{O}(n)$-time seeded extraction with $O(\log(n/\varepsilon))$ seed length and large output length for all ranges of parameters.
@@ 1.4 Future Work
  • Remove the preprocessing step that our constructions behind ˜1 require when $k$ or $\varepsilon$ are small. We expand on some promising approaches suggested by Jesse Goodman and also some barriers we face in section˜3.3.3.
  • On the practical side, develop software implementations of seeded extractors with near-optimal seed length and large output length. In particular, we think that our non-recursive construction in section˜4.1 holds promise in this direction.
@@ 1.5 Acknowledgements
We thank Jesse Goodman for many valuable comments and suggestions that greatly improved this work, and in particular for suggesting the approach outlined in section˜3.3.3.
@@ 2.1 Notation
We often use uppercase Roman letters to denote sets and random variables – the distinction will be clear from context. We denote the support of a random variable $X$ by $\mathsf{supp}(X)$, and for a random variable $X$ and set $S$, we also write $X\sim S$ to mean that $X$ is supported on $S$. For a random variable $X$, we write $x\sim X$ to mean that $x$ is sampled according to the distribution of $X$. We use $U_{d}$ to denote a random variable that is uniformly distributed over $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$. For two strings $x$ and $y$, we may write $(x,y)$ for their concatenation. Given two random variables $X$ and $Y$, we denote their product distribution by $X\times Y$ (i.e., $\Pr[X\times Y=(x,y)]=\Pr[X=x]\cdot\Pr[Y=y]$). Given a positive integer $n$, we write $n]=\{1,\dots,n\}$. For a prime power $q$, we denote the finite field of order $q$ by $\mathds{F}_{q}$. We denote the base-$2$ logarithm by $\log$.
@@ 2.2 Model of Computation
We work in the standard, multi-tape, Turing machine model with some fixed number of work tapes. In particular, there exists a constant $C$ such that all our claimed time bounds hold whenever we work with at most $C$ work tapes. This also implies that our results hold in the RAM model, wherein each machine word can store integers up to some fixed length, and standard word operations take constant time. In section˜5 we will give, in addition to the standard Turing machine model bounds, an improved runtime bound that is dedicated to the logarithmic-cost RAM model.
2.3 Fast Finite Field Operations
Understanding the complexity of operations in finite fields will be useful for the analysis of the complexity of the condensers from [GUV09, KT22] in section˜3.3. For a prime power $q=p^{\ell}$, we let $M_{q}(d)$ be the number of field operations required to multiply two univariate polynomials over $\mathds{F}_{q}$ of degree less than $d$, and $M_{q}^{\mathsf{b}}(d)$ be the bit complexity of such a multiplication, so $M_{q}^{\mathsf{b}}(d)\leq M_{q}(d)\cdot T(q)$, where we denote by $T(q)$ an upper bound on the bit complexity of arithmetic operations in $\mathds{F}_{q}$. Harvey and van der Hoeven [Hv19] (see also [Hv22]) showed that
$M^{\mathsf{b}}_{q}(d)=O(d\log q\cdot\log(d\log q)\cdot 4^{\log^{*}(d\log q)}).$
Overall, when $d\leq q$, we have that $M_{q}^{\mathsf{b}}(d)=d\log d\cdot\widetilde{O}(\log q)$. [^11] A similar bound can be obtained using simpler methods. If $\mathds{F}_{q}$ contains a $d$-th root of unity, one can get $M_{q}(d)=d\log d$ from the classic FFT algorithm [21]. For a simpler algorithm attaining the bound $M_{q}(d)=d\log d\operatorname{loglog}d$, see [vG13, Sections 8, 10]. When $p=2$, $M_{q}(d)=d\log d\operatorname{loglog}d$ also follows from Schönhage’s algorithm [SCH77]. Now, since $M_{q}^{\mathsf{b}}(d)=M_{q}(d)\cdot M_{p}(\ell)\cdot M_{p}^{\mathsf{b}}(0)$, a bound of $d\log d\cdot\widetilde{O}(\log q)$ also follows.
We will use fast multi-point evaluation and fast computation of derivatives (together with the preceding bounds on $M_{q}^{\mathsf{b}}$).
Lemma 2.1 ([BM74], see also [vG13, Chapter 10]).
Let $d\in\mathbb{N}$, and let $q=p^{r}$ be a prime power. Then, given a polynomial $f\in\mathds{F}_{q}[X]$ of degree at most $d$ (together with a representation of $\mathds{F}_{q}$ via an irreducible polynomial over $\mathds{F}_{p}$ of degree $r$), the following holds.
@@ 2.3 Fast Finite Field Operations
    For $t\leq d$ and $\alpha\in\mathds{F}_{q}$, one can compute the derivatives $f(\alpha),f^{\prime}(\alpha),\ldots,f^{(t)}(\alpha)$ in time $O(M_{q}^{\mathsf{b}}(d)\cdot\log d)=d\log^{2}d\cdot\widetilde{O}(\log q)$.
Note that when $q\leq 2^{d}$, we can bound $O(M_{q}(d)\cdot\log d)$ by $\widetilde{O}(d)\cdot\log q$.
For a comprehensive discussion of fast polynomial arithmetic, see Von Zur Gathen and Gerhard’s book [vG13] (and the more recent important developments [Hv21]).
@@ 2.4 Statistical Distance, Entropy
We present some relevant definitions and lemmas about statistical distance and min-entropy.
@@
$\Delta(X,Y)=\max_{\mathcal{T}\subseteq\mathcal{S}}|\Pr[X\in\mathcal{T}]-\Pr[Y\in\mathcal{T}]|=\frac{1}{2}\sum_{x\in\mathcal{S}}|\Pr[X=x]-\Pr[Y=x]|.$
@@
The min-entropy rate of $X$ is given by $\frac{\mathbf{H}_{\infty}(X)}{\log|\mathcal{X}|}$.
@@ 2.5 Extractors and Condensers
A random variable $X$ is an $((n_{1},n_{2},\dots,n_{t}),(k_{1},k_{2},\dots,k_{t}))$-block source if we can write $X=(X_{1},X_{2},\dots,X_{t})$, each $X_{i}\in\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n_{i}}$, where $\widetilde{\mathbf{H}}_{\infty}(X_{i}|X_{1},\ldots,X_{i-1})\geq k_{i}$ for all $i\in[s]$. In the special case where $k_{i}=\alpha n_{i}$ for all $i\in[t]$, we say that $X$ is an $((n_{1},n_{2},\dots,n_{t}),\alpha)$-block source.
We say that $X$ is an exact block source if $\mathbf{H}_{\infty}(X_{i}|X_{1}=x_{1},\ldots,X_{i-1}=x_{i-1})\geq k_{i}$ for any prefix $x_{1},\dots,x_{i-1}$. lemma˜2.6 tells us that any $((n_{1},n_{2},\ldots,n_{t}),\alpha)$-block-source is $\varepsilon$-close to an exact $((n_{1},n_{2},\ldots,n_{t}),(1-\zeta)\alpha)$-block-source, where $\varepsilon=\sum_{i=1}^{t}2^{-\alpha\zeta n_{i}}$.
@@
where $Y$ is uniformly distributed over $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$ and is independent of $X$ and $U_{m}$ is uniformly distributed over $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m}$. We say that $\mathsf{Ext}$ is strong if $(\mathsf{Ext}(X,Y),Y)\approx_{\varepsilon}U_{m+d}$.
@@
$(\mathsf{Ext}(X,Y),Y,W)\approx_{\varepsilon}(U_{m+d},W),$
where $Y$ is uniformly distributed over $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$ and is independent of $X$, and $U_{m+d}$ is uniformly distributed over $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m+d}$ and independent of $W$.
@@
By lemma˜2.6, every strong $(k,\varepsilon)$-seeded extractor $\mathsf{Ext}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\to\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m}$ is also an average-case strong $(k^{\prime}=k+\log(1/\varepsilon),\varepsilon^{\prime}=2\varepsilon)$-seeded extractor.
@@
A function $\mathsf{Cond}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\rightarrow\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m}$ is a $(k,k^{\prime},\varepsilon)$ (seeded) condenser if the following holds. For every $(n,k)$-source $X$, it holds that $Z=\mathsf{Cond}(X,Y)$ is $\varepsilon$-close to some $Z^{\prime}$ with $\mathbf{H}_{\infty}(Z^{\prime})\geq k^{\prime}$, where $Y$ is uniformly distributed over $\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}$ and is independent of $X$.
We say that $\mathsf{Cond}$ is strong if $(Y,\mathsf{Cond}(X,Y))$ is $\varepsilon$-close to some distribution $(Y,Z)$ with min-entropy $k^{\prime}$ (and note that here, necessarily, $d$ bits of entropy come from the seed). Finally, we say that $\mathsf{Cond}$ is lossless if $k^{\prime}=k+d$.
We also define seeded extractors tailored to block sources.
@@
A function $\mathsf{BExt}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n_{1}}\times\cdots\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n_{t}}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\to\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{m}$ is a $(k_{1},\dots,k_{t},\varepsilon)$ strong block-source extractor if for any $((n_{1},\dots,n_{t}),(k_{1},\dots,k_{t}))$-block-source $X$,
$(\mathsf{BExt}(X,Y),Y)\approx_{\varepsilon}U_{m+d},$
@@
We will also require the following extractors based on the leftover hash lemma and fast hash functions. We state a result from [TSS+11] which requires seed length $d\approx 2m$, where $m$ is the output length.
Lemma 2.13 (fast hash-based extractors [TSS+11, Theorem 10], adapted. See also [HT16, Table I]).
@@
Note that by appending the seed to the output of the extractor, we can get the following: There exists a constant $c$ such that for any constant $\theta\leq\frac{1}{3}$, $d\geq c\log(n/\varepsilon)$ and $k\geq\theta d+c\log(1/\varepsilon)$, there exists a $(k,\varepsilon)$-seeded extractor $\mathsf{Ext}\colon\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{n}\times\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{d}\rightarrow\mathopen{}\mathclose{{\left\{{0,1}}}\right\}^{(1+\theta)d}$.